If mass of one atom is 3.32x 10^-23 g, then calculate number of nucleons (neutrons and protons) present in 2 atoms of the element :
1
Expert's answer
2020-11-12T09:32:09-0500

Since the mass of protons and neutrons is equal, the number of nucleons in two atoms

"N=2\times \frac{m_a}{m_p}=2\times \frac{3.32\times 10^{-23}}{1.67\times 10^{-27}}\approx 39760" .